OpenStudy (anonymous):

Help please?? <3 Find S12 for 1 + 6 + 11 + 16 +… 308 322 336 342

hartnn (hartnn):

so you know the formula for Sum of terms in Arithmetic series?

OpenStudy (anonymous):

no

hartnn (hartnn):

ok, here you go! \(\Large S_n = n/2 (2a+ (n-1)d) \) where a = First term = 1 d = common difference = .... ? can you find? and n = number of terms = 12, because we need sum of 12 terms

OpenStudy (anonymous):

how do you find the common difference?

hartnn (hartnn):

d = common difference = difference between next term and current term.. like d = 2nd term - 1st term = 3rd term - 2nd term = ..and so on now can you find the d?

OpenStudy (anonymous):

so the common difference is 5

hartnn (hartnn):

thats right! d= 5 :) now plug in all the values in the formula :)

OpenStudy (anonymous):

\[s _{12}=12/2(2\times1+(12-1)5)\] Right?

hartnn (hartnn):

yup, go on!

OpenStudy (anonymous):

\[s _{12}=12/2(2\times1(11)5)\]

hartnn (hartnn):

so whats your final answer ?

OpenStudy (anonymous):

i got 12/114 but that's not an option so... I don't know what I did wrong?

hartnn (hartnn):

12/2 is separate which will evaluate to 6 6 (2+ 55) = ... ?

OpenStudy (anonymous):

oh okay so it 342

hartnn (hartnn):

yes! thats correct :) the whole \(2a + (n-1)d\) is in numerator and not in denominator :)